A 1,200 kg car ascends a road inclined at 200 to the horizontal at 25 km/h. Neglecting frictional opposition, estimate the power required to keep going at this speed.


1
Expert's answer
2021-11-14T17:11:50-0500

power required to keep going:

"P=m\vec{g}\cdot \vec{v}=mgvsin\alpha=1200\cdot 9.8\cdot25sin20\degree/3.6=27932" W